# Algorithmic stability implies training-conditional coverage for distribution-free prediction methods

### v2

@@ 1 Introduction
To make this goal more precise, define the training data set [^1] Here and throughout the paper, we will write “data set” to denote a collection of data points, even though the same data point may potentially appear multiple times—that is, more formally, we might refer to $\mathcal{D}_{n}$ as a multiset of data points. as $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1,\dots,n}$, and define the conditional noncoverage rate as
@@
1.1 Prior results.
Prior work has considered the question of training-conditional coverage for a range of commonly used distribution-free predictive inference methods: split conformal prediction, full conformal prediction, and jackknife+ (we will give background on each of these methods shortly). Vovk (2012) establishes a training-conditional coverage bound of the type (3) for split conformal prediction. In contrast, Bian and Barber (2023) prove that for the full conformal prediction and jackknife+ methods, no such bound can be established without further assumptions; their work constructs specific counterexamples where the worst-case scenario outlined in (2) can occur for these two methods. Moreover, as demonstrated empirically by Bian and Barber (2023, Section 5), both the jackknife+ and full conformal method can fail to have training-conditional coverage even in a simple linear regression setting, if the sample size and dimension are approximately equal (in which case the problem becomes ill-conditioned, leading to instability). Hence, Bian and Barber (2023) conjecture that an assumption of algorithmic stability may be sufficient to ensure training-conditional coverage for these methods.
@@ 2.1 Distribution-free prediction
With the problem setting in place, we now review three methods for distribution-free prediction—split conformal, full conformal, and the jackknife+. (For further background on these methods, including questions of how the choice of the base algorithm $\mathcal{A}$ affects the resulting prediction interval in terms of length or other properties, and how the resulting prediction intervals compare with more traditional model-based methods, see, e.g., Vovk et al. (2005); Lei et al. (2018); Sesia and Candès (2020).)
@@ 2.1.2 Full conformal prediction
is the fitted model when the algorithm $\mathcal{A}$ is run on the training data together with hypothesized test point $(X_{n+1},y)$. (We refer to $\widehat{C}^{\textnormal{CP}}_{n}(X_{n+1})$ as a prediction set, rather than prediction interval, since in principle the construction may return a set that is not a connected interval; in practice, however, it is often the case that $\widehat{C}^{\textnormal{CP}}_{n}(X_{n+1})$ is indeed an interval.)
@@
i.e., the example essentially attains the worst-case scenario described in (2). (Bian and Barber (2023, Section 5) also demonstrate empirically that full conformal can fail to achieve training-conditional coverage even with a simple linear regression algorithm, in an ill-conditioned regime.)
@@ 2.2 Algorithmic stability
Some algorithms, such as ridge regression and bagging (Bühlmann and Yu, 2002; Grandvalet, 2004; Elisseeff et al., 2005; Soloff et al., 2024a, b), are known to satisfy algorithmic stability properties due to their construction. However, most modern modeling algorithms are too complex to be analyzed theoretically (i.e., to determine whether a stability property is guaranteed to hold). Recent work by Kim and Barber (2023) establishes that testing stability empirically is in fact impossible—any test that provides an assumption-free “certificate” of stability cannot have power that is much better than random.
Another line of work that connects stability and inference is closely related to the field of differential privacy, which is typically referred to as adaptive data analysis. There, a notion of stability is extracted from differential privacy and is used to construct valid inferences when hypotheses are determined in a data-driven way (Dwork et al., 2015a, b; Rogers et al., 2016). Zrnic and Jordan (2023) build on this idea to derive a framework for post-selection inference using stability achieved by randomization. We note that the notions of stability these works considered are different from ours, and typically much stronger: the privacy literature typically considers uniform notions of stability, i.e., assumptions of the flavor $\max_{i\in[n]}\left|\widehat{\mu}(X_{n+1})-\widehat{\mu}_{-i}(X_{n+1})\right|=\mathrm{o}_{P}(1)$ (which requires that no training point is allowed to be overly influential, while weaker notions of stability only require that a randomly selected training point should not be overly influential).
@@ 3.1 Results for jackknife+
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma$-J+}}(\mathcal{D}_{n})<\alpha+3\sqrt{\frac{\log(1/\delta)}{2\min\{n,m\}}}+2\sqrt[3]{\frac{2\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)}{\gamma}}\right\}\geq 1-3\delta-\sqrt[3]{\frac{2\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)}{\gamma}},$
@@ 3.2 Results for full conformal
$\widehat{C}^{\textnormal{$\gamma$-CP}}_{n}(X_{n+1})=\left\{y\in\mathbb{R}:|y-\widehat{\mu}^{y}(X_{n+1})|\leq\textnormal{Quantile}\big{(}\{|Y_{i}-\widehat{\mu}^{y}(X_{i})|\}_{i\in[n]};(1-\alpha)(1+1/n)\big{)}+\gamma\right\}.$
@@
We might ask whether, for full conformal, we can also prove a result for the uninflated prediction set (as in Theorem 3.3, for jackknife+). However, without further assumptions, we cannot derive this type of corollary for full conformal. This is because for jackknife+, the prediction set is in fact an interval, and so inflated set $\widehat{C}_{n}^{\textnormal{$\gamma$-J+}}(X_{n+1})$ and its uninflated counterpart $\widehat{C}_{n}^{\textnormal{J+}}(X_{n+1})$ differ only by Lebegue measure $2\gamma$—that is, $\textnormal{Leb}(\widehat{C}_{n}^{\textnormal{$\gamma$-J+}}(X_{n+1})\backslash\widehat{C}_{n}^{\textnormal{J+}}(X_{n+1}))=2\gamma$. For full conformal, since there is no equivalent simple characterization of how the sets $\widehat{C}_{n}^{\textnormal{CP}}(X_{n+1})$ and $\widehat{C}_{n}^{\textnormal{$\gamma$-CP}}(X_{n+1})$ differ, the measure of the set difference $\textnormal{Leb}(\widehat{C}_{n}^{\textnormal{$\gamma$-CP}}(X_{n+1})\backslash\widehat{C}_{n}^{\textnormal{CP}}(X_{n+1}))$ could potentially be arbitrarily large.
@@ 3.3 An asymptotic view
(For full conformal, we can also ask the analogous question with in-sample stability in place of out-of-sample stability.)
We will now see that, despite appearances, our theorems above actually ensure a guarantee of this seemingly stronger type—that is, asymptotically, $1$-stability (with a vanishing value of $\beta^{\textnormal{out}}_{1,n}(\mathcal{A},P)$ or $\beta^{\textnormal{in}}_{1,n}(\mathcal{A},P)$) ensures training-conditional coverage.
@@
Fix any algorithm $\mathcal{A}$ and distribution $P$. Let $\gamma_{n}$ be any positive and nonincreasing sequence such that $\beta_{1,n-1}^{\textnormal{out}}(\mathcal{A},P)/\gamma_{n}\rightarrow 0$. Then there exist sequences $\epsilon_{n},\delta_{n}\rightarrow 0$ (constructed explicitly in the proof) such that
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma_{n}$-J+}}(\mathcal{D}_{n})\leq\alpha+\epsilon_{n}\right\}\geq 1-\delta_{n}$
for all $n\geq 1$. In other words, $\alpha_{P}^{\textnormal{$\gamma_{n}$-J+}}(\mathcal{D}_{n})\leq\alpha+\mathrm{o}_{p}(1)$.
Moreover, if the conditional distribution of $Y\mid X$ has density $f_{Y|X}(y\mid x)$ with respect to Lebesgue measure, with $\mathbb{E}\left[\sup_{y\in\mathbb{R}}f_{Y|X}(y\mid X)\right]<\infty$, then the same result holds for the uninflated jackknife+ interval: if $\beta_{1,n}^{\textnormal{out}}(\mathcal{A},P)\to 0$, then there exist sequences $\epsilon_{n},\delta_{n}\rightarrow 0$ (constructed explicitly in the proof) such that
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})\leq\alpha+\epsilon_{n}\right\}\geq 1-\delta_{n}$
for all $n\geq 1$. In other words, under this additional assumption, $\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})\leq\alpha+\mathrm{o}_{p}(1)$.
@@
Fix any algorithm $\mathcal{A}$ and distribution $P$. Let $\gamma_{n}$ be any positive and nonincreasing sequence such that $\beta_{1,n+1}^{\textnormal{in}}(\mathcal{A},P)/\gamma_{n}\rightarrow 0$. Then there exist sequences $\epsilon_{n},\delta_{n}\rightarrow 0$ (constructed explicitly in the proof) such that
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma_{n}$-CP}}(\mathcal{D}_{n})\leq\alpha+\epsilon_{n}\right\}\geq 1-\delta_{n}$
for all $n\geq 1$. In other words, $\alpha_{P}^{\textnormal{$\gamma_{n}$-CP}}(\mathcal{D}_{n})\leq\alpha+\mathrm{o}_{p}(1)$.
These results show us that, as long as $1$-stability holds with some vanishing parameter $\beta_{1,n}(\mathcal{A},P)$ (i.e., out-of-sample or in-sample stability, as appropriate, for jackknife+ or for full conformal), then we can always choose an inflation value $\gamma_{n}$ (with $\gamma_{n}$ tending to zero slower than $\beta_{1,n}(\mathcal{A},P)$) such that asymptotic training-conditional coverage will hold (or, if we additionally assume $Y\mid X$ has a continuous distribution, then for jackknife+ no inflation is necessary). Thus we can conclude that, from this asymptotic view, $1$-stability is indeed sufficient for establishing training-conditional coverage. To give a preview of the proofs, the idea is that we take a sequence $m_{n}\rightarrow\infty$ that grows sufficiently slowly, and relate $m_{n}$-stability (which is needed for applying the finite-sample theorems) to $1$-stability (which is assumed in the asymptotic theorems).
@@ 4.2 Stability when data points are swapped rather than removed/added
$\widehat{\mu}_{n}=\mathcal{A}\left((X_{1},Y_{1}),\dots,(X_{n},Y_{n})\right),\ \ \widehat{\mu}_{n}^{\prime}=\mathcal{A}\left((X_{1},Y_{1}),\dots,(X_{n-m},Y_{n-m}),(X^{\prime}_{n-m+1},Y^{\prime}_{n-m+1}),\dots,(X^{\prime}_{n},Y^{\prime}_{n})\right),$
@@
$\bar{\beta}_{m,n}^{\textnormal{out}}(\mathcal{A},P)\leq 2\beta_{m,n}^{\textnormal{out}}(\mathcal{A},P)$
@@
Let $\widehat{\mu}_{n+m}=\mathcal{A}((X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X^{\prime}_{n-m+1},Y^{\prime}_{n-m+1}),\dots,(X^{\prime}_{n},Y^{\prime}_{n}))$. For out-of-sample stability, we have
@@
by definition of out-of-sample $m$-stability (together with the assumption of symmetry of $\mathcal{A}$). A similar argument holds for in-sample stability as well. ∎
@@
The finite-sample training-conditional coverage guarantees given in Theorems 3.2 and 3.3 for the jackknife+, hold with $\bar{\beta}_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)$ in place of $\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)$. Similarly, the asymptotic guarantees given in Theorem 3.6 hold as well, under the modified assumption $\bar{\beta}_{1,n-1}^{\textnormal{out}}(\mathcal{A},P)/\gamma_{n}\rightarrow 0$ (or correspondingly $\bar{\beta}_{1,n}^{\textnormal{out}}(\mathcal{A},P)\rightarrow 0$ for uninflated jackknife+ if $Y|X$ has a continuous distribution).
@@
The finite-sample training-conditional coverage guarantee given in Theorem 3.5 for full conformal, holds with $\bar{\beta}_{m-1,n+1}^{\textnormal{in}}(\mathcal{A},P)$ in place of $\beta_{m-1,n+1}^{\textnormal{in}}(\mathcal{A},P)$. Similarly, the asymptotic guarantee given in Theorem 3.7 holds as well, under the modified assumption $\bar{\beta}_{1,n+1}^{\textnormal{in}}(\mathcal{A},P)/\gamma_{n}\rightarrow 0$.
For jackknife+, we will see that this swap-stability version of our main result holds because we can establish a weaker sort of converse to Proposition 4.3, for out-of-sample stability. (For full conformal, the proof will require a different approach.)
Before stating the partial converse result, we need to introduce some notation. Recalling that an algorithm $\mathcal{A}$ maps a data set of any size to a fitted function, we will write $\mathcal{A}_{n}$ to denote the map defined by applying $\mathcal{A}$ to a training set of size $n$. That is, $\mathcal{A}_{n}$ is a map with input space $(\mathcal{X}\times\mathbb{R})^{n}$, and $\mathcal{A}$ is simply the collection of maps $\mathcal{A}_{n}$ across all $n$.
Proposition 4.6.
Proof of Proposition 4.6.
Now we have
Proof of Theorem 4.4 (jackknife+ with swap-stability).
First, we can observe that $\widehat{C}_{n}^{\textnormal{$\gamma$-J+}}(X_{n+1})$ depends on the algorithm $\mathcal{A}$ only through the map $\mathcal{A}_{n-1}$; that is, when running jackknife+, $\mathcal{A}$ is only ever trained on data sets of size $n-1$. Now let $\mathcal{A}^{\prime}$ be defined as in Proposition 4.6 (except with $n-1$ in place of $n$), so that we have $\mathcal{A}^{\prime}_{n-1}=\mathcal{A}_{n-1}$. In other words, at sample size $n-1$, running jackknife+ with $\mathcal{A}$ is exactly equivalent to running jackknife+ with algorithm $\mathcal{A}^{\prime}$. Therefore, the results of Theorems 3.2 and 3.3 can be applied with $\mathcal{A}^{\prime}$ in place of $\mathcal{A}$; applying Proposition 4.6 tells us that $\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A}^{\prime},P)\leq\bar{\beta}_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)$, which verifies the desired finite-sample training-coverage guarantee. Finally, the asymptotic guarantees with swap-stability are proved exactly as for (non-swap) stability, i.e., following the same arguments as in the proof of Theorem 3.6. ∎
4.3 In-sample, out-of-sample, or neither?

### v3

@@ 1 Introduction
To make this goal more precise, define the training data set111 Here and throughout the paper, we will write “data set” to denote a collection of data points, even though the same data point may potentially appear multiple times—that is, more formally, we might refer to $\mathcal{D}_{n}$ as a multiset of data points. as $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1,\dots,n}$, and define the conditional noncoverage rate as
@@
1.1 Prior results
Prior work has considered the question of training-conditional coverage for a range of commonly used distribution-free predictive inference methods: split conformal prediction, full conformal prediction, and jackknife+. We will give background on each of these methods shortly in Section 2. Vovk (2012) establishes a training-conditional coverage bound of the type (3) for split conformal prediction. In contrast, Bian and Barber (2023) prove that for the full conformal prediction and jackknife+ methods, no such bound can be established without further assumptions; their work constructs specific counterexamples where the worst-case scenario outlined in (2) can occur for these two methods. Moreover, as demonstrated empirically by Bian and Barber (2023, Section 5), both the jackknife+ and full conformal method can fail to have training-conditional coverage even in a simple linear regression setting, if the sample size and dimension are approximately equal (in which case the problem becomes ill-conditioned, leading to instability). Hence, Bian and Barber (2023) conjecture that an assumption of algorithmic stability may be sufficient to ensure training-conditional coverage for these methods.
@@ 2.1 Distribution-free prediction
With the problem setting in place, we now briefly review three methods for distribution-free prediction—split conformal, full conformal, and the jackknife+. For further background on these methods, including questions of how the choice of the base algorithm $\mathcal{A}$ affects the resulting prediction interval in terms of length or other properties, and how the resulting prediction intervals compare with more traditional model-based methods, see, e.g., Vovk et al. (2005); Lei et al. (2018); Sesia and Candès (2020).
@@ 2.1.2 Full conformal prediction
is the fitted model when the algorithm $\mathcal{A}$ is run on the training data together with hypothesized test point $(X_{n+1},y)$. Here, we refer to $\widehat{C}^{\textnormal{CP}}_{n}(X_{n+1})$ as a prediction set, rather than prediction interval, since in principle the construction may return a set that is not a connected interval; in practice, however, it is often the case that $\widehat{C}^{\textnormal{CP}}_{n}(X_{n+1})$ is indeed an interval.
@@
i.e., the example essentially attains the worst-case scenario described in (2). Bian and Barber (2023, Section 5) also demonstrate empirically that full conformal can fail to achieve the training-conditional coverage even with a simple linear regression algorithm, in an ill-conditioned regime.
@@ 2.2 Algorithmic stability
Some algorithms, such as ridge regression and bagging (Bühlmann and Yu, 2002; Grandvalet, 2004; Elisseeff et al., 2005; Soloff et al., 2024a, b), are known to satisfy algorithmic stability properties due to their construction. However, most modern modeling algorithms are too complex to be analyzed theoretically, i.e., determining whether a stability property is guaranteed to hold is difficult. Recent work by Kim and Barber (2023) establishes that testing stability empirically is in fact impossible—any test that provides an assumption-free “certificate” of stability cannot have power that is much better than random.
Another line of work that connects stability and inference is closely related to the field of differential privacy, which is typically referred to as adaptive data analysis. There, a notion of stability is extracted from differential privacy and is used to construct valid inferences when hypotheses are determined in a data-driven way (Dwork et al., 2015a, b; Rogers et al., 2016). Zrnic and Jordan (2023) build on this idea to derive a framework for post-selection inference using stability achieved by randomization. We note that the notions of stability these works considered are different from ours, and typically much stronger: the privacy literature typically considers uniform notions of stability, i.e., assumptions of the flavor $\max_{i\in[n]}\left|\widehat{\mu}(X_{n+1})-\widehat{\mu}_{-i}(X_{n+1})\right|=\mathrm{o}_{P}(1)$, which requires that no training point is allowed to be overly influential, while weaker notions of stability only require that a randomly selected training point should not be overly influential.
@@ 3.1 Results for jackknife+
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma$-J+}}(\mathcal{D}_{n})<\alpha+3\sqrt{\frac{\log(1/\delta)}{2\min\{n,m\}}}+2\sqrt[3]{\frac{2\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)}{\gamma}}\right\}\geq 1-3\delta-\sqrt[3]{\frac{2\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)}{\gamma}}$
@@ 3.2 Results for full conformal
$\widehat{C}^{\textnormal{$\gamma$-CP}}_{n}(X_{n+1})=\Big{\{}y\in\mathbb{R}:|y-\widehat{\mu}^{y}(X_{n+1})|\leq\textnormal{Quantile}\big{(}\{|Y_{i}-\widehat{\mu}^{y}(X_{i})|\}_{i\in[n]};(1-\alpha)(1+1/n)\big{)}+\gamma\Big{\}}.$
@@
We might ask whether, for full conformal, we can also prove a result for the uninflated prediction set (as in Theorem 3.3, for jackknife+, which assumes only a bounded density for $Y\mid X$). However, it is likely that a bounded density assumption is not sufficient for deriving an analogous result for full conformal. This is because for jackknife+, the prediction set is in fact an interval, and so inflated set $\widehat{C}_{n}^{\textnormal{$\gamma$-J+}}(X_{n+1})$ and its uninflated counterpart $\widehat{C}_{n}^{\textnormal{J+}}(X_{n+1})$ differ only by Lebegue measure $2\gamma$—that is, $\textnormal{Leb}(\widehat{C}_{n}^{\textnormal{$\gamma$-J+}}(X_{n+1})\backslash\widehat{C}_{n}^{\textnormal{J+}}(X_{n+1}))=2\gamma$. For full conformal, since there is no equivalent simple characterization of how the sets $\widehat{C}_{n}^{\textnormal{CP}}(X_{n+1})$ and $\widehat{C}_{n}^{\textnormal{$\gamma$-CP}}(X_{n+1})$ differ, the measure of the set difference $\textnormal{Leb}(\widehat{C}_{n}^{\textnormal{$\gamma$-CP}}(X_{n+1})\backslash\widehat{C}_{n}^{\textnormal{CP}}(X_{n+1}))$ could potentially be arbitrarily large.
@@ 3.3 An asymptotic view
For full conformal, we can also ask the analogous question with in-sample stability in place of out-of-sample stability.
We will now see that our theoretical guarantees that rely on $m$-stability, can actually ensure a guarantee of this seemingly stronger type—that is, asymptotically, training-conditional coverage requires only $1$-stability (with a vanishing value of $\beta^{\textnormal{out}}_{1,n}(\mathcal{A},P)$ or $\beta^{\textnormal{in}}_{1,n}(\mathcal{A},P)$). To ease notation in the theorem statements, we define $\beta^{*,\textnormal{out}}_{1,n}=\sup_{n^{\prime}\geq n}\beta^{\textnormal{out}}_{1,n^{\prime}}$, and similarly $\beta^{*,\textnormal{in}}_{1,n}=\sup_{n^{\prime}\geq n}\beta^{\textnormal{in}}_{1,n^{\prime}}$. Since, in general, we expect stability to improve with increasing sample size $n$, in practical settings we would likely have $\beta^{*,\textnormal{out}}_{1,n}={\beta}^{\textnormal{out}}_{1,n}$, and same for in-sample stability.
@@
Fix any sequences $b_{n},\gamma_{n}>0$ such that $b_{n}/\gamma_{n}\to 0$. Then there exists a sequence $\epsilon_{n}\rightarrow 0$ (constructed explicitly in the proof) such that
$\lim\limits_{n\to\infty}\sup\limits_{\begin{subarray}{c}\mathcal{A},P\textnormal{ s.t. }\\ \beta_{1,n-1}^{*,\textnormal{out}}(\mathcal{A},P)\leq b_{n}\end{subarray}}\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma_{n}$-J+}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}=0.$
Moreover, a similar result holds without inflation, under an additional bounded density assumption. For each $B>0$, define a set $\mathcal{P}_{B}$ of distributions on $(X,Y)\in\mathcal{X}\times\mathbb{R}$, given by
[…]
Fix any sequences $b_{n},B_{n}>0$ such that $b_{n}B_{n}\to 0$. Then there exists a sequence $\epsilon_{n}\rightarrow 0$ (constructed explicitly in the proof) such that
$\lim\limits_{n\to\infty}\sup\limits_{\begin{subarray}{c}\mathcal{A},P\textnormal{ s.t. $P\in\mathcal{P}_{B_{n}}$,}\\ \beta_{1,n-1}^{*,\textnormal{out}}(\mathcal{A},P)\leq b_{n}\end{subarray}}\mathbb{P}\left\{\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}=0.$
@@
Fix any sequences $b_{n},\gamma_{n}>0$ such that $b_{n}/\gamma_{n}\to 0$. Then there exists a sequence $\epsilon_{n}\rightarrow 0$ (constructed explicitly in the proof) such that
$\lim\limits_{n\to\infty}\sup\limits_{\begin{subarray}{c}\mathcal{A},P\textnormal{ s.t. }\\ \beta_{1,n+1}^{*,\textnormal{in}}(\mathcal{A},P)\leq b_{n}\end{subarray}}\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma_{n}$-CP}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}=0.$
To help parse these asymptotic results, consider fixing the training algorithm $\mathcal{A}$ and data distribution $P$. Then, these results show us that, as long as $1$-stability holds with some vanishing parameter $\beta_{1,n}(\mathcal{A},P)$ (i.e., out-of-sample or in-sample stability, as appropriate, for jackknife+ or for full conformal), then we can always choose an inflation value $\gamma_{n}$, with $\gamma_{n}\to 0$ sufficiently slowly, such that asymptotic training-conditional coverage will hold (or, if we additionally assume $Y\mid X$ has a continuous distribution, then for jackknife+ no inflation is necessary). Thus we can conclude that, from this asymptotic view, $1$-stability is indeed sufficient for establishing training-conditional coverage. To give a preview of the proofs, the idea is that we take a sequence $m_{n}\rightarrow\infty$ that grows sufficiently slowly, and relate $m_{n}$-stability (which is needed for applying the finite-sample theorems) to $1$-stability (which is assumed in the asymptotic theorems).
@@ 4.2 Stability when data points are swapped rather than removed/added
$\bar{\beta}_{m,n}^{\textnormal{out}}(\mathcal{A},P)\leq 2\beta_{m,n}^{\textnormal{out}}(\mathcal{A},P),$
@@
Let
$\widehat{\mu}_{n+m}=\mathcal{A}((X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X^{\prime}_{n-m+1},Y^{\prime}_{n-m+1}),\dots,(X^{\prime}_{n},Y^{\prime}_{n})).$
For out-of-sample stability, we have
@@
by definition of out-of-sample $m$-stability together with the assumption of symmetry of $\mathcal{A}$. A similar argument holds for in-sample stability as well. ∎
@@
The finite-sample training-conditional coverage guarantees given in Theorems 3.2 and 3.3 for the jackknife+, hold with $\bar{\beta}_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)$ in place of $\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)$. Similarly, the asymptotic guarantees given in Theorem 3.6 hold as well, with $\bar{\beta}^{*,\textnormal{out}}_{1,n-1}=\sup_{n^{\prime}\geq n-1}\bar{\beta}^{\textnormal{out}}_{1,n^{\prime}}$ in place of $\beta^{*,\textnormal{out}}_{1,n-1}$.
@@
The finite-sample training-conditional coverage guarantee given in Theorem 3.5 for full conformal, holds with $\bar{\beta}_{m-1,n+1}^{\textnormal{in}}(\mathcal{A},P)$ in place of $\beta_{m-1,n+1}^{\textnormal{in}}(\mathcal{A},P)$. Similarly, the asymptotic guarantee given in Theorem 3.7 holds as well, with $\bar{\beta}^{*,\textnormal{in}}_{1,n+1}=\sup_{n^{\prime}\geq n+1}\bar{\beta}^{\textnormal{in}}_{1,n^{\prime}}$ in place of $\beta^{*,\textnormal{in}}_{1,n+1}$.
For jackknife+, this swap-stability version of our main result holds because we can establish a weaker sort of converse to Proposition 4.3, for out-of-sample stability. For full conformal, the proof will require a different approach. See the Appendix B for details.
4.3 In-sample, out-of-sample, or neither